8. Which of the following is the equation in slope-intercept form of the line through the points (1, 4) and (–1, –7)? (2 points)
Answer by rfer(16322)

m=-7-4/-1-1
m=-11/-2
m=5.5
-------------------
y=mx+b
-7=5.5*-1+b
-7=-5.5+b
b=-1.5
----------------
y=5.5x-1.5